The management of Kunkel Company is considering the purchase of a $36,000 machine that would reduce operating costs by $8,500 per year. At the end of the machine’s five-year useful life, it will have zero scrap value. The company’s required rate of return is 13%.

Use Excel or a financial calculator to solve.

Required:
1a.

Determine the net present value of the investment in the machine. (Any cash outflows should be indicated by a minus sign. Round answers to the nearest dollar.)

Net present value $(6,106)
2b.

What is the difference between the total, undiscounted cash inflows and cash outflows over the entire life of the machine? (Any cash outflows should be indicated by a minus sign.)

Item Cash Flow Years Total Cash Flows
Annual cost savings $1 5 $5
Initial investment $1 1 1
Net cash flow $4
Answer #1

1.purchase of machine -36000

scrap value    0

cost savings *annuity factor    8500*3.517=29894.5

so net present value is -36000+29894.5=-6105.5 or 6106

2.difference between the total undiscounted cash flows

= -36000+(8500*5)=-36000+42500=6500

positive net present value of 6500.